# GARG-AML against Smurfing: A Scalable and Interpretable Graph-Based Framework for Anti-Money Laundering

### v2

@@ 3.2 Undirected
Given the sights of the example above, we define the general GARG-AML method, starting with the adjacency matrix. The first row of the adjacency matrix, which corresponds to entries $(A_{1,j})$, represents the connections of node $v$ under consideration. The next m1m-1 rows, i.e., $(A_{i,j})$ for $2\leq i\leq m$}, represent the connections of nodes that are located at a distance of two from $v$. The final $n$ rows, i.e., $(A_{i,j})$ for $m+1\leq i\leq m+n$}, represent the connections of nodes at a distance of one from $v$.
Enforcing this node ordering—the target node, second-order neighbours, and first-order neighbours—for the adjacency matrix ensures that the diagonal blocks remain sparse (low density), and the off-diagonal blocks have a high density in case we are faced with a pure smurfing pattern. The exact order of nodes within each group is not important. For example, any permutation within the first-order neighbours, such as $C,B,D$, is equally valid.
We consider three primary blocks of the adjacency matrix for score calculation: the upper-left, upper-right, and lower-right blocks. These are defined as follows:
@@
The density, i.e., the relative number of non-zero elements in each block are essential for our method. GARG-AML measures the similarity between an observed pattern and the pure smurfing template by analysing the expected density and actual density (or sparsity) of each block. This results in a single quantitative score expressing similarity to the pure smurfing pattern.
Several structural properties of the adjacency matrix are taken into account to come to more rigorous density calculations. First, self-loops are excluded; thus, diagonal elements are always zero and are not considered from the sparsity/density calculations. Furthermore, in $\text{block}_{1}$ (upper-left block), the first row and column can never contain ones, as a ‘1’ would indicate a direct connection between a node and its neighbours at distance two—a contradiction. This upper-left block serves as a check that the nodes at distance two do no directly make transactions to each other, in case multiple nodes at a distance of two are present. Similarly, for $\text{block}_{2}$ (upper-right block), all the elements in the first row must be ones since these represent connections between the node under investigation and its first-order neighbours.
We consider only the free entries, i.e., those that are not certain upfront, for density calculations. For readability, we define the block size as the number of free elements. This is reflected in both the numerator (number of non-zero elements) and denominator (block size).
@@
$\text{score}_{2}$ $=$ $\frac{\text{sum}(\text{block}_{2})-m}{mn-n}$ (6)
@@
Smurfing is likely to be present when the density of $\text{block}_{2}$ (upper-right block) is high, while $\text{block}_{1}$ (upper-left block) and $\text{block}_{3}$ (lower-right) exhibit densities close to zero. Therefore, the final GARG-AML score is calculated as=:
@@
where $l_{i}$ is a weight given to each term in the score calculation. The inclusion of $l_{i}$ increases the flexibility of the score calculation. If $l_{i}=1$, the score is the average of $\text{score}_{1}$ and $\text{score}_{3}$. In our experiments, however, we take a weighted average using the block sizes:
$l_{1}$ $=$ $m^{2}-3m+2$
$l_{3}$ $=$ $n^{2}-n.$
@@
The final score in Equation (8) ranges from 1-1 to $1$. When the second-order neighbourhood resembles a smurfing pattern, the final score is close to 1. To illustrate how these calculations are done, additional examples are provided in Appendix A.1.
3.3 Directed
A defining trait of smurfing is the uni-directional flow of funds. Money moves from one account to others with minimal or no return flow (Starnini et al., 2021; Lee et al., 2020; Li et al., 2020; Shadrooh & Nørvåg, 2024), forming what we call pure smurfing patterns (see Figure 3). To capture this behaviour, we extend GARG-AML to directed networks.
Figure 3: The directed version of the smurfing pattern example, extended with an additional sending node ($A^{\prime}$) and receiving node ($E^{\prime}$). Nodes at level 0 are put in blue, at level 1 in purple and at level 2 in green.
We, again, start with an example. In the directed network of Figure 3, unlike the undirected case (Figure 2), node $A$ only sends funds to the smurfs, while node $E$ only receives from them. This leads to a non-symmetric adjacency matrix, requiring the analysis of nine matrix blocks instead of three (see Equation (9)). To ensure consistency in block-wise density calculations—especially with multiple senders or receivers—nodes at distance two are classified as senders (level 0) and receivers (level 2). When considering the network on the right in Figure 3, we want that node $A^{\prime}$ is put at the same level as node $A$ and node $E^{\prime}$ at the same level as node $E$. The resulting adjacency matrix is then
Accordingly, we define two types of second-order neighbourhoods in network $G(V,E)$ for the general GARG-AML score calculations. The strong second-order neighbourhood, denoted $\mathcal{N}^{s}_{2}(v)$, considers nodes reachable from $v$ via a directed path of length two. The weak second-order neighbourhood, $\mathcal{N}^{w}_{2}(v)$, is based on the undirected network. In both, only nodes at exact distance two are included.
@@ 3.3 Directed
To account for reverse flow patterns we repeat this selection on the reversed network. This means that a node $w$ a distance two of node $v$ is put at level 0 if there is no directed path of length two from $v$ to $w$ nor from $w$ to $v$.
As in the undirected case, structural constraints of the adjacency matrix are considered when calculating block-wise densities. For the upper-left block, $\text{score}_{00}$, the first row and column are excluded since all other level-0 nodes are nodes at distance two. Additionally, diagonal elements in all on-diagonal blocks ($\text{score}_{00}$, $\text{score}_{11}$ and $\text{score}_{22}$) are excluded from density calculations, as self-loops are not present.
Adjustments are also made to $\text{score}_{20}$ and $\text{score}_{02}$, which represent connections between the node and its second-order neighbours. The first row of $\text{block}_{02}$ (from the node to its second-order neighbours) and first column of $\text{block}_{20}$ (from the second-order neighbours to the node) are always zero and are therefore excluded.
@@
We take the sizes to be $\mid\mathcal{N}_{1}(v)\mid=n$, $\mid\mathcal{N}^{s}_{2}(v)\mid=m$, and $\mid\mathcal{N}^{w}_{2}(v)\setminus\mathcal{N}^{s}_{2}(v)\mid=l-1$, resulting in $l$ nodes at level 0, $n$ nodes at level 1, and $m$ nodes at level 2. The full adjacency matrix is partitioned into nine blocks, as shown in Equation (11), with block labels using zero-based indexing to reflect node levels 0, 1, and 2.
@@
To compute the final score, we take the weighted average of the two blocks expected to have high density — according to the smurfing pattern in Equation (9) — and subtract the weighted average of the remaining blocks, which are expected to be sparse. The use of the average ensures that the final GARG-AML score is between -1 and 1, to have consistent comparison and interpretation across datasets.
@@
To illustrate how these calculations are done, a small example is provided in Appendix A.2.
The GARG-AML scores offer three key advantages. First, the scores are close to 1 when a node’s second-order neighbourhood resembles a smurfing pattern, providing a clear and interpretable signal for both automated detection and human investigation. Second, the scores can easy be integrated into existing workflows. The densities and final score can be seen as additional feature values that can be added when training a model or setting up business rules. Third, the method is highly scalable. Each score is computed independently using only a node’s second-order neighbourhood, allowing for efficient parallel computation and low memory usage, even on large transaction networks.
@@ 4.1 Data
We experimentally evaluate the detection power of GARG-AML on multiple datasets. A first set concerns synthetic networks that we generate ourselves using various existing methods, to illustrate how GARG-AML scales with growing transaction networks and to evaluate how well it operates when faced with camouflaging tactics. In addition, the performance of GARG-AML is tested on two realistic open-source datasets. These mimic real-life transaction data; they include millions of observations of transactions, with only very few of these confirmed money laundering cases.
@@ 4.1.1 Synthetic Data
We generate synthetic networks of varying sizes using three generation methods, i.e., Barabási-Albert (Barabási & Albert, 1999), Erdős-Rényi (Erdős & Rényi, 1959) and Watts–Strogatz (Watts & Strogatz, 1998). The resulting networks have a range of different properties, e.g., scale-free, small world, etc., that allow the evaluation of GARG-AML across different network topologies.
Each of these methods has specific parameters that we vary to obtain a range of networks to test our method across different conditions. We summarize the ranges of these parameters in Table 3. The first parameter is the size of the network, which is subsequently set to 100, 10,000 and 100,000.
@@
The other parameters have slightly different definitions depending on the method. The number of edges is set at 1, 2 or 5. For Barabási-Albert, this represents the number of nodes to which each new node connects, resulting in a large difference in degree across the nodes. For Watts–Strogatz, on the other hand, this is half the initial degree that all the nodes have before some of the edges are rewired.
The edge (rewiring) probability is set to 0.001 and 0.01. The Erdős-Rényi method uses this parameter to express the probability of having an edge between any two nodes. The Watts–Strogatz method starts with a fixed number of nodes, and uses this parameter to express the probability that an edge changes its end node (rewiring).
Smurfing patterns are injected after network generation. The three types of smurfing patterns that we consider in the experiments, and which are visualized in Figure 4, are the following:
@@
Figure 4: Illustration of a random network in which three different methods of injecting smurfing patterns are presented. The solid lines indicate nodes and edges of the original network, whereas the dotted lines indicate injected nodes and edges that correspond to smurfing.
In our experiments, each of the smurfing patterns is included either three or five times; see Table 3. The number of smurfs in each pattern is randomly chosen between two and ten. This is done to allow for smaller as well as larger smurfing patterns. A summary of all the synthetic datasets obtained is presented in Table 4.
Table 4: Summary of the numbers of nodes and edges and money laundering percentages for the synthetic datasets according to the considered network generation parameters.
@@
Since we have information on which nodes are involved in which pattern for each dataset, we not only test the methods for money laundering detection in general; we also analyse how the methods perform on separate patterns. Similarly, we investigate how performance changes when money launderers apply higher levels of obfuscation methods.
@@ 4.1.2 Open-Source Data
As post-hoc injection of specific patterns may lead to detection bias (Deprez et al., 2024) and altered topology (Egressy et al., 2024), we also include data without this injection in the experiments to evaluate GARG-AML.
We use the HI-Small and LI-Large datasets, made publicly available by Altman et al. (2023) and Egressy et al. (2024), representing a relatively small dataset with high money laundering activity and a large dataset with low money laundering activity, respectively. Table 5 summarises the number of nodes, edges and the percentage of cases labelled as money laundering. Especially LI-Large closely resembles real-world transaction networks.
@@
The two datasets are generated using the same simulator (Altman et al., 2023; Egressy et al., 2024), which models financial transactions in a virtual world involving individuals, companies, and banks. A subset of these transactions is labelled as money laundering. Since the full transaction graph is available, the authors were able to assign detailed labels, categorizing transaction streams into eight distinct types: fan-out, fan-in, gather-scatter, scatter-gather, simple cycle, random, bipartite, and stack.
Among these, the scatter-gather and gather-scatter patterns—illustrated in Figure 1—represent the primary smurfing behaviour that GARG-AML aims to detect. The distribution of labelled patterns is shown in Table 6. It is worth noting that some transactions are labelled as money laundering but do not match any of the eight defined structural patterns. These are grouped under the category “not classified”, indicating that they are suspicious but not attributable to a specific laundering archetype.
Table 6: The distribution of the different types of money laundering transaction patterns for the two synthetic datasets.
@@
The original labels in these datasets are provided at the edge level, while GARG-AML produces features at the node level. To bridge this gap, we aggregate labels per client by computing the ratio of labelled laundering transactions to the total number of transactions associated with each node. This ratio is interpreted as a propensity score for money laundering.
To generate binary labels, we compare each node’s propensity score against a set of predefined thresholds: $10\%,20\%,30\%,50\%$ and $90\%$. A node is labelled as involved in money laundering if the proportion of its laundering-labelled transactions exceeds the selected threshold. The resulting label distributions are presented in Table 7 and Table 8, corresponding to the HI-Small and LI-Large datasets, respectively.
@@
Table 7: Percentage of instances labelled as 1 in the HI-Small dataset, for the different patterns at different thresholds.
@@
Table 8: Percentage ($\%$) of instances labelled as 1 in the LI-Large dataset, for the different patterns at different thresholds.
@@ 4.2 Model Specifications
The full GARG-AML calculation pipeline consists of multiple stages, beginning with pre-processing steps that prepare the network data for analysis, followed by the computation of the GARG-AML scores. We also introduce a possible extension of the GARG-AML method that applies tree-based methods. We describe this use cases in detail below to illustrate the versatility of GARG-AML. An overview of the complete pipeline is shown in Figure 5 and summarised in Algorithm 1.
@@
Transaction networks often exhibit a scale-free structure (Starnini et al., 2021), which are characterized by the presence of a few nodes (accounts) with disproportionately high numbers of incoming and outgoing transactions. As a result, second-order neighbourhoods can become very large and can encompass an unmanageable number of nodes for score computation.
@@
Therefore, we first partition the network into communities using the Louvain method (Blondel et al., 2008). This method consistently shows state-of-the-art performance for community detection, and scales well to large graphs (Lancichinetti & Fortunato, 2009; Yang et al., 2016). Communities are identified on the basis of the undirected version of the transaction network, and only intra-community edges are retained—edges between communities are discarded.
@@
However, the Louvain method can still produce overly large communities (Fortunato & Barthélemy, 2007; Newman & Girvan, 2004). To mitigate this, we use the implementation from NetworkX (Hagberg et al., 2008), which includes a resolution parameter inspired by the improvements proposed by Traag et al. (2019). In our experiments, we set this parameter to 10, which effectively encourages the algorithm to form smaller, more manageable communities, thus maintaining tractability in the GARG-AML score computation.
@@
After pre-processing the network using the Louvain method, we calculate the block densities as described above via Equation (8) and Equation (3.3).
As a possible extension, we apply a decision tree and gradient boosting classifier with features on the basis of the GARG-AML scores. These include the original score as before, which is enhanced with summary statistics of each node’s first-order neighbours, i.e., the minimal, mean and maximal values of the GARG-AML scores as well as the standard deviation of the neighbours’ scores. Since successful smurfing involves many smurfs, we also include each node’s degree and the summary statistics—minimum, mean, maximum and standard deviation—of the degrees of the node’s neighbours.
The goal of including these extra features is to capture more complex neighbourhood structures that may not be evident from the scores alone. In particular, this helps reduce false positives—for example, a linear chain of three nodes may yield a perfect score, even though a single “smurf” is insufficient evidence of an actual laundering operation.
@@
Applying this on the undirected and directed network results in four additional model. We use a 70-30 stratified train-test split for evaluation. This split is performed on the resulting feature table, not the network itself. For the decision tree, we use the Gini criterion and require that each leaf contain at least ten samples. The gradient boosting classifier consists of 100 trees with a maximal depth of three and a learning rate of 0.1. We require that each leaf node should contain at least 10 observations.
We benchmark GARG-AML and its extension against established state-of-the-art methods, namely, AutoAudit (Lee et al., 2020) and FlowScope (Li et al., 2020). For both methods, we use the implementations provided by the authors and retain their default settings.
@@
Various small adaptations are necessary. AutoAudit returns an optimal reordering of indices in the adjacency matrix. However, this reordering affects only part of the matrix and may depend on the arbitrary labelling of nodes. To ensure robustness, we repeat the experiments three times with random initialization of the node indices and report the average results. Both methods also return an ordering of transactions resulting from the $n$ different iterations. To reflect the decreasing level of certainty, the first group gets prediction $1$, the second $\frac{n-1}{n}$ and so on until the final, $n$th, group gets score $\frac{1}{n}$. All other nodes that were not returned by the methods get score 0.

### v3

@@ 3.2 Undirected Network
Given these insights, we define the general GARG-AML method, starting with the adjacency matrix. The first row of the adjacency matrix, which corresponds to entries $(A_{1,j})$, represents the connections of node $v$ under consideration. The next m1m-1 rows, i.e., $(A_{i,j})$ for $2\leq i\leq m$}, represent the connections of nodes at distance two from $v$. The final $n$ rows, i.e., $(A_{i,j})$ for $m+1\leq i\leq m+n$}, represent the connections of nodes at distance one.
Enforcing this node ordering—the target node, second-order neighbours, and first-order neighbours—for the adjacency matrix ensures that the diagonal blocks remain sparse (low density), and the off-diagonal blocks have a high density for pure smurfing patterns. The exact order of nodes within each group is not important. Any permutation within the first-order neighbours, such as $C,B,D$, is equally valid.
We consider three blocks of the adjacency matrix for score calculation: the upper-left, upper-right, and lower-right blocks. These are defined as follows:
@@
Our method uses the density, i.e., the relative number of non-zero elements in each block. GARG-AML measures the similarity between an observed pattern and the pure smurfing template by analysing the expected density and actual density (or sparsity) of each block. This results in a single quantitative score expressing similarity to the pure smurfing pattern.
Several structural properties of the adjacency matrix are taken into account for more rigorous density calculations. First, self-loops are excluded; thus, diagonal elements are always zero. Furthermore, in $\text{block}_{1}$ (upper-left block), the first row and column can never contain ones, as a ‘1’ would indicate a direct connection between a node and its neighbours at distance two—a contradiction. This upper-left block serves as a check that the nodes at distance two do not make transactions to each other. Similarly, for $\text{block}_{2}$ (upper-right block), all the elements in the first row must be ones since these represent connections between the node under investigation and its first-order neighbours.
We consider only the free entries, i.e., those that are not certain upfront, for density calculations. We define the block size as the number of free elements. This is reflected in the numerator (number of non-zero elements) and denominator (block size).
@@
$\text{score}_{2}$ $=$ $\frac{\text{sum}(\text{block}_{2})-n}{mn-n}$ (6)
@@
Smurfing is likely to be present when the density of $\text{block}_{2}$ (upper-right block) is high, while $\text{block}_{1}$ (upper-left block) and $\text{block}_{3}$ (lower-right) exhibit low densities. Therefore, the final GARG-AML score is calculated as:
@@
where $l_{i}$ is a weight given to each term in the score calculation. If $l_{i}=1$, the score is the average of $\text{score}_{1}$ and $\text{score}_{3}$. We take a weighted average using the block sizes:
$l_{1}$ $=$ $m^{2}-3m+2$ (9)
$l_{3}$ $=$ $n^{2}-n.$ (10)
@@
The final score in Equation (8) ranges from 1-1 to $1$. Second-order neighbourhood resembling a smurfing pattern result in a score close to 1. Additional examples are provided in Appendix A.1 as illustration.
3.3 Directed Network
A defining trait of smurfing is the uni-directional flow of funds. Money moves from one account to others with minimal/no return flow (Starnini et al., 2021; Lee et al., 2020; Li et al., 2020; Shadrooh and Nørvåg, 2024), forming pure smurfing patterns (see Figure 3). To capture this behaviour, we extend GARG-AML to directed networks.
Figure 3: The directed version of the smurfing pattern example, extended with an additional sending node (A’) and receiving node (E’). Nodes at level 0 are put in blue, at level 1 in purple and at level 2 in green.
We start with an example. In the directed network of Figure 3 node $A$ only sends funds to the smurfs, while node $E$ only receives from them. This leads to a non-symmetric adjacency matrix, requiring the analysis of nine matrix blocks instead of three (see Equation (12)). To ensure consistency in block-wise density calculations—especially with multiple senders or receivers—nodes at distance two are classified as senders (level 0) and receivers (level 2).
We define two types of second-order neighbourhoods in network $G(V,E)$. The strong second-order neighbourhood, denoted $\mathcal{N}^{s}_{2}(v)$, considers nodes reachable from $v$ via a directed path of length two. The weak second-order neighbourhood, $\mathcal{N}^{w}_{2}(v)$, is based on the undirected network. In both, only nodes at exact distance two are included.
@@ 3.3 Directed Network
To account for reverse flow patterns, we repeat this selection on the reversed network. This means that a node $w$ a distance two of node $v$ is put at level 0 if there is no directed path of length two from $v$ to $w$ nor from $w$ to $v$. The other second-order neighbours are put at level 2.
For the extended network in Figure 3, node $A^{\prime}$ is put at the same level as node $A$ and node $E^{\prime}$ at the same level as node $E$. The resulting adjacency matrix is then
Structural constraints of the adjacency matrix are considered when calculating block-wise densities. For the upper-left block, $\text{score}_{00}$, the first row and column are excluded since all other level-0 nodes are nodes at distance two. Additionally, diagonal elements are excluded for $\text{score}_{00}$, $\text{score}_{11}$ and $\text{score}_{22}$, as self-loops are not present.
Adjustments are also made to $\text{score}_{20}$ and $\text{score}_{02}$, which represent connections between the node and its second-order neighbours. The first row of $\text{block}_{02}$ (from the node to its second-order neighbours) and first column of $\text{block}_{20}$ (from the second-order neighbours to the node) are always zero.
@@
We take the sizes to be $\mid\mathcal{N}_{1}(v)\mid=n$, $\mid\mathcal{N}^{s}_{2}(v)\mid=m$, and $\mid\mathcal{N}^{w}_{2}(v)\setminus\mathcal{N}^{s}_{2}(v)\mid=l-1$, resulting in $l$ nodes at level 0, $n$ nodes at level 1, and $m$ nodes at level 2. The adjacency matrix is partitioned into nine blocks, as shown in Equation (13), with block labels using zero-based indexing to reflect node levels 0, 1, and 2.
@@
The final score takes the weighted average of the two blocks expected to have high density—according to the smurfing pattern in Equation (12)—and subtract the weighted average of the remaining blocks, which are expected to be sparse. The final GARG-AML score is between -1 and 1, to have consistent comparison and interpretation across datasets.
@@
Additional examples are provided in Appendix A.2 as illustration.
The GARG-AML scores offer three key advantages. First, scores near 1 offer a transparent signal for automated detection and human investigation by reflecting structural smurfing resemblances. Second, the resulting scores and densities can be easily incorporated into existing business rules or machine learning workflows. Third, each score is computed independently using a node’s second-order neighbourhood, allowing for efficient parallel computation and low memory usage, even on large transaction networks.
3.4 Integration in Existing Detection Systems
This section illustrates how GARG-AML can easily be integrated in existing detection methods.
Rule-based systems can integrate new scenarios based on the GARG-AML score and its interaction with other features. The rules remain highly interpretable, given that the GARG-AML scores are within a fixed range, with higher scores indicating that the network resembles smurfing. More complex scenarios can be constructed using the individual densities and sizes of the blocks. These are obtained during the intermediary calculations, so these do not increase calculation costs.
Detection models based on machine learning can also be extended. The scores or densities can be interpreted as additional client features. The existing pipelines can be kept to retrain the models.
@@ 4.1 Data
Due to the high sensitivity of financial transaction and fraud data, which precludes the use of proprietary banking datasets, we evaluate the detection power of GARG-AML using a two-tiered approach. First, we utilize self-generated networks created through various simulation methods to maintain full control over network size, topology, and the specific smurfing patterns present, allowing us to illustrate the model’s scalability and its performance against diverse camouflaging tactics. Second, we test the framework on realistic open-source datasets that mimic real-life transaction flows—including millions of transactions with a realistically low ratio of confirmed money laundering cases–to establish a robust evaluation procedure while ensuring all experiments remain fully reproducible.
@@ 4.1.1 Synthetic Data
We generate synthetic networks of varying sizes using three generation methods, i.e., Barabási-Albert (BarabásiAlbert-László1999EoSi), Erdős-Rényi (Erdős and Rényi, 1959) and Watts–Strogatz (Watts and Strogatz, 1998). By generating networks with varying properties, we evaluated GARG-AML across a spectrum of topologies.
Each method has specific parameters that we vary to obtain a range of networks to test our method. We summarize these in Table 3. The first parameter is the size of the network, which is subsequently set to 100, 10,000 and 100,000.
@@
The definition of the other parameters differ depending on the method. For Barabási-Albert, the number of edges represents the number of nodes to which each new node connects, resulting in a large difference in degree across the nodes. For Watts–Strogatz, on the other hand, this is half the initial degree that all the nodes have before some of the edges are rewired.
For the Erdős-Rényi method, the edge (rewiring) probability is the probability of having an edge between any two nodes. The Watts–Strogatz method starts with a fixed number of nodes, and uses this parameter to express the probability that an edge changes its end node (rewiring).
Smurfing patterns are injected after network generation. The three types of smurfing patterns (visualized in Figure 4) considered in the experiment are:
@@
Figure 4: A random network with the three methods of smurfing pattern injection. Solid lines indicate nodes and edges of the original network. Dotted lines indicate injected nodes and edges.
In our experiments, each of the smurfing patterns is included either three or five times; see Table 3. The number of smurfs in each pattern is randomly chosen between two and ten. This allows for both smaller and larger smurfing patterns. A summary of all synthetic datasets is presented in Table 4.
Table 4: Summary of the numbers of nodes and edges and money laundering percentages for the synthetic datasets according to the network generation parameters.
@@
Since we have information on which nodes are involved in which pattern, we not only test the methods for money laundering detection in general; we analyse the performance on the different patterns separately, representing increasing levels of obfuscation.
@@ 4.1.2 Open-Source Data
As post-hoc injection of specific patterns may lead to detection bias (Deprez et al., 2024) and altered topology (Egressy et al., 2024), we also include data without this injection. We use the HI-Small and LI-Large datasets, made publicly available by Altman et al. (2023) and Egressy et al. (2024), representing a relatively small dataset with high money laundering activity and a large dataset with low money laundering activity, respectively. Table 5 summarises the number of nodes, edges and the percentage of cases labelled as money laundering. Especially LI-Large closely resembles real-world transaction networks.
@@
These two datasets are generated using the same simulator (Altman et al., 2023; Egressy et al., 2024). It models financial transactions in a virtual world involving individuals, companies, and banks. A subset of these transactions is labelled as money laundering. The data contains detailed labels, categorizing transaction streams into eight distinct types: fan-out, fan-in, gather-scatter, scatter-gather, simple cycle, random, bipartite, and stack. The distribution of labelled patterns is shown in Table 6.
It is worth noting that some transactions are labelled as money laundering but do not match any of the eight defined structural patterns. These are grouped under the category “not classified”, indicating that they are suspicious but not attributable to a specific laundering archetype.
Table 6: The distribution of the money laundering transaction patterns for the two synthetic datasets.
@@
The original labels in these datasets are provided at the edge level, while GARG-AML produces features at the node level. To bridge this gap, we aggregate labels by computing the ratio of labelled laundering transactions to the total number of transactions per client. This ratio is interpreted as a propensity score for money laundering.
We compare each node’s propensity score against a set of predefined cut-offs: $10\%,20\%,30\%,50\%$ and $90\%$. A node is labelled as money launderer if the proportion of its laundering-labelled transactions exceeds said cut-off. The resulting label distributions are presented in Table 7 and Table 8, for the HI-Small and LI-Large datasets, respectively.
@@
Table 7: Percentage of instances labelled as 1 in the HI-Small dataset per cut-offs.
@@
Table 8: Percentage ($\%$) of instances labelled as 1 in the LI-Large dataset per cut-off.
@@ 4.2 Model Specifications
The full GARG-AML calculation pipeline consists of multiple stages, beginning with pre-processing steps that prepare the network data for analysis, followed by the computation of the GARG-AML scores. We introduce a possible extension that applies tree-based methods since these have been shown to have strong performance for AML (Jullum et al., 2020; Ramadhan, 2024; Halford et al., 2025; Lokanan, 2025). We describe this use cases in detail below to illustrate the versatility of GARG-AML. An overview of the complete pipeline is shown in Figure 5 and summarised in Algorithm 1.
@@
Transaction networks often exhibit a scale-free structure (Starnini et al., 2021), which is characterized by a majority of nodes having only few connections, while a few nodes, called hubs, have a disproportionately high numbers of connections. As a result, second-order neighbourhoods that contain such hubs encompass an unmanageable number of nodes for score computation.
@@
Therefore, we first partition the network into communities using the Louvain method (Blondel et al., 2008). This method consistently shows state-of-the-art performance for community detection, and scales well to large graphs (Lancichinetti and Fortunato, 2009; Yang et al., 2016). Communities are identified on the basis of the undirected version of the transaction network, and only intra-community edges are retained—edges between communities are discarded.
@@
The Louvain method can produce overly large communities (Fortunato and Barthélemy, 2007; Newman and Girvan, 2004). To mitigate this, we use the implementation from NetworkX (Hagberg et al., 2008), which includes a resolution parameter inspired by the improvements proposed by Traag et al. (2019). In our experiments, we set this parameter to 10 to obtain smaller, more manageable communities.
@@
After pre-processing, we calculate the block densities as described above. We come to the final undirected and directed score via Equation (8) and Equation(3.3), respectively.
As a possible extension, we apply a decision tree and gradient boosting classifier that use features based on the GARG-AML scores. A decision tree is a non-parametric model that performs greedy binary partitioning of the features on an impurity criterion. The resulting model can be seen as a collection of IF-ELSE rules. The gradient boosting classifier is an ensemble learning method that constructs an additive model in a forward stage-wise manner. It optimizes a differentiable loss function by sequentially training ‘weak learners’ (shallow decision trees), where each subsequent tree is fit to the negative gradient of the loss function (the residuals) of the previous ensemble.
To capture the coordinated nature of smurfing, the base GARG-AML scores were supplemented with neighbourhood summary statistics (min, mean, max, and standard deviation of neighbour degrees and scores). This structural context helps the model differentiate between suspicious multi-mule topologies and benign linear chains, significantly reducing false positives.
@@
Applying this on the undirected and directed network results in four additional model. We use a 70-30 stratified train-test split for evaluation. This split is performed on the resulting feature table, not the network itself. For the decision tree, we use the Gini criterion and each leaf should contain at least ten samples. The gradient boosting classifier consists of 100 trees with a maximal depth of three and a learning rate of 0.1. Each leaf node should contain at least 10 observations.
@@
We benchmark GARG-AML and its extension against established state-of-the-art methods, namely, AutoAudit (Lee et al., 2020) and FlowScope (Li et al., 2020), with their default hyperparameters. Various small adaptations are necessary. AutoAudit returns an optimal reordering of indices in the adjacency matrix. This affects only part of the matrix and may depend on the arbitrary numbering of nodes. To ensure robustness, we repeat the experiments three times with random initialization of the node indices and report the average results.
Both methods return an ordering of transactions over $n$ iterations. To reflect the decreasing level of certainty, the first group gets prediction $1$, the second $\frac{n-1}{n}$ and so on until the final, $n$th, group gets score $\frac{1}{n}$. All other nodes get score 0.